OpenStudy (anonymous):

Find the equation of the line tangent to sin(xy)+y^4= 1 at (0,1).

hartnn (hartnn):

know implicit differentiation ?

OpenStudy (anonymous):

yes, have to solve for y prime. Im a little confused on what to do with the sin(xy), product rule or chain rule?

hartnn (hartnn):

chain rule first, because its the function of function of x. then u need product rule for xy, try it

OpenStudy (anonymous):

for the chain rule, would it be cos(xy) * y ' ?

hartnn (hartnn):

should i give one next step ?

hartnn (hartnn):

d/dx (sin(xy)) = cos (xy) d/dx(xy) this is chain rule, got it ?

hartnn (hartnn):

d/dx(xy) should be simple product rule.....

OpenStudy (anonymous):

y cos (xy) ?

hartnn (hartnn):

did u take y as constant and so (xy)' is just y ? NO, y is the function of x, so u need product rule like this : (xy)' = x'y + xy' now what is x' = ?

hartnn (hartnn):

got that^ ?

OpenStudy (anonymous):

lol, so sorry im just drawing blanks right now for some reason!

hartnn (hartnn):

so should i give you solution so that u can look up or you want to try on your own first ?

OpenStudy (anonymous):

hmmmm solution so I can see how its done if you dont mind.

hartnn (hartnn):

sin(xy)+y^4= 1 diff. w.r.t x (sin(xy)+y^4)'= 0 (sin(xy))'+(y^4)'= 0 cos(xy) (xy)' + 4y^3 y' =0 cos (xy) (x'y+y'x) +4y^3 y' =0 now x'=d/dx(x)=1 cos (xy) (y+y'x) +4y^3 y' =0 put point (0,1), that is put x=0,y=1 cos 0 = 1 (1+0) + 4 y' = 0 y' = -1/4 = slope of tangent at point (0,1) clear till here ? ask if there is any doubt in any step.

OpenStudy (anonymous):

cos(xy)(y+y'x)+4y^3y'=0 after that step I'm a little confused. how do you start to solve for y'.

hartnn (hartnn):

the slope of tangent at a point (x1,y1) is given by putting x=x1, y= y1 in y'= d/dx(f(x)) so i put y=1, x=0 in cos(xy)(y+y'x)+4y^3y'=0

OpenStudy (anonymous):

okay, sorry it took so long but i see it now lol.

hartnn (hartnn):

so now u have slope of line, there is a point on that line (0,1) can u find the equation of line ?

OpenStudy (anonymous):

one question first, could you factor out y' to help somewhat simplify?

OpenStudy (anonymous):

like in cos(xy)(y+y'x)+4y^3*y'

hartnn (hartnn):

oh u just want y' from cos(xy)(y+y'x)+4y^3*y' =0 ,right ? y cos (xy) + y'(xcos xy) + y' (4y^3) = 0 can u find y' from here ?

OpenStudy (anonymous):

xcosxy +4y^3 / ycos(xy) ?

hartnn (hartnn):

y cos (xy) + y'(xcos xy) + y' (4y^3) = 0 y'(xcos xy) + y' (4y^3) = - y cos (xy) y' = - y cos (xy) / (x cos xy +4y^3)

hartnn (hartnn):

if u now put x=0, y=1, u still get y'=slope = -1/4

OpenStudy (anonymous):

so would the tangent line be y-1=(-1/4)(x-0)

hartnn (hartnn):

absolutely! just simplify it ...

OpenStudy (anonymous):

okay so y=-1/4x +1 ?

hartnn (hartnn):

that is in slope intercept form,correct. is it mentioned in which form the equation is needed ?

OpenStudy (anonymous):

no but our professor used it in that form, because we have to approximate after solving for y.

hartnn (hartnn):

then keep it in that form. so u got all steps ?

OpenStudy (anonymous):

yes lol. thank you so much, and sorry it took so long!

hartnn (hartnn):

no problem . welcome ^_^ have a nice day :)

OpenStudy (anonymous):

you too! :)